# Talent Hoarding in Organizations

### v2

@@
Abstract
Most organizations rely on managers to identify talented workers. However, managers who are evaluated on team performance have an incentive to hoard workers. This study provides the first empirical evidence of talent hoarding using personnel records and survey evidence from a large manufacturing firm. Talent hoarding is reported by three-fourths of managers, is detectable in managerial decisions, and occurs more frequently when hoarding incentives are stronger. Using quasi-random variation in exposure to talent hoarding, I demonstrate that hoarding deters workers from applying to new positions, inhibiting worker career progression and altering the allocation of talent in the firm.
@@ 1 Introduction
Firms must continually decide how to allocate workers to jobs, a process which has critical implications for productivity (Rosen, 1982, Holmstrom and Tirole, 1989). Because it is difficult to perfectly observe worker ability, most firms rely on managers to identify talented workers who can be promoted to higher-level positions. However, when a talented worker leaves their team for a new position, team performance suffers. Since managers are rewarded based on team performance and firms cannot perfectly monitor manager actions, the conflicting interests of manager and firm create the potential for moral hazard (Holmstrom, 1979). A growing body of evidence documents that workers in high-level positions have large impacts on firm performance (Bloom and Van Reenen, 2007, Lazear et al., 2015), implying that managers may create significant efficiency costs if they hoard talented workers rather than promote them. Ample anecdotal and survey evidence points to widespread talent hoarding in organizations. In a global survey, half of organizations report that managers hoard talent by discouraging worker mobility (i4cp, 2016). A US-based survey finds that workers in one-third of firms feel the need to keep internal applications secret from their managers out of fear of retaliation (KornFerry, 2015). In Germany, 83% of the top publicly listed companies cite managerial talent hoarding as a key friction in their organization (hkp, 2021). [^2] Talent hoarding also occurs in science and academia. Zuckerman (2021) documents how Katalin Karikó, a seminal developer of mRNA vaccines and 2023 Nobel prize winner, experienced talent hoarding when she decided to leave her lab for a new position. Her advisor Robert Suhadolnik vowed “to do whatever he could to stop his protégée from leaving…he made it clear she had two career choices. ‘You can work in my lab or go home,’ he told her. Suhadolnik followed through on this threat, telling a local immigration office that she was living in the country illegally and should be deported.” Despite the apparent prevalence of talent hoarding and its likely detrimental consequences, very little empirical evidence on talent hoarding exists in economics. Studying talent hoarding empirically is challenging. Managers often hoard talent through hidden actions that are difficult to observe, even in rich datasets. Furthermore, identifying the causal impacts of talent hoarding requires plausibly exogenous variation in hoarding. This study provides the first empirical evidence on talent hoarding in organizations. I combine rich personnel records from a large manufacturing firm with survey evidence from both managers and workers in the firm. I document widespread managerial incentives to hoard talent: the majority of managers perceive a conflict of interest with respect to developing talent, because more developed workers are more likely to leave the team. Talent hoarding is reported by 75% of managers. Using a direct measure of talent hoarding derived from the firm’s personnel records, I test predictions from my conceptual framework and confirm that stronger hoarding incentives lead to more talent hoarding. I leverage quasi-random variation in workers’ exposure to talent hoarding to demonstrate the impacts of talent hoarding on workers’ career outcomes. Hoarding significantly deters workers’ applications for internal positions, many of which would have been successful. Because talent hoarding has a larger impact on high-quality workers, hoarding reduces the quality of the applicant pool, thus restricting the firm’s ability to effectively allocate talent within the organization.
I develop a simple conceptual framework that formalizes the misaligned incentives that give rise to talent hoarding. Since managers are both tasked with identifying productive workers and are rewarded based on team productivity, managers have an incentive to hoard talent by preventing workers from seeking new positions outside the team. This framework yields empirical predictions for both manager and worker behavior. For example, when gains to hoarding are larger or worker departure from the team is costlier, managers engage in more hoarding. High-productivity workers who experience more hoarding are more deterred from applying, which negatively impacts internal applicant pools and has the potential to distort the allocation of talent within the firm. To empirically test for talent hoarding, I collect a unique combination of personnel records, internal job application data, and new survey evidence from a large manufacturing firm that employs over 200,000 workers. I focus my analysis on the firm’s largest internal labor market, consisting of over 30,000 white-collar and management employees in Germany. I demonstrate that the firm is similar to other large firms in terms of its workforce and organizational design. As in many other large firms, employees who want to switch to a new position or to be promoted are required to apply for the position. Since most workers in the same team work in similar positions, workers who seek different positions, including promotions, typically need to apply to jobs outside the team. The data I assemble offer three key advantages. First, collecting new survey evidence from managers in the firm allows me to directly document that managers face misaligned incentives and act accordingly. Second, the firm’s personnel records allow me to construct a measure of talent hoarding using the extent to which managers systematically suppress the visibility of workers. Without such data, directly measuring managerial talent hoarding is empirically challenging because hoarding by definition involves hidden actions. Third, by combining personnel records with the universe of application and hiring decisions at the firm, I am able to document the extent to which talent hoarding deters workers’ high-stakes applications that would otherwise have been successful. Survey responses from 62% of managers in the firm reveal that incentives to hoard talent are widespread. The majority of managers (55%) explicitly acknowledge that supporting workers’ career development entails a conflict of interest because more developed workers are more likely to leave the team. This conflict of interest relates to the misaligned incentives managers face: while almost all managers (96%) believe they would have a large impact on workers’ careers, only 36% perceive this impact to be valued by the firm. A further disincentive to develop talent is the risk of workers leaving the team: 68% of managers state that they would be more likely to support workers’ career development if it would be easier to replace lost talent. The survey demonstrates that managerial talent hoarding actions, which are reported by 75% of managers, are a common response to these prevailing managerial incentives. Managers’ revealed actions provide further evidence that managers purposefully hoard talent. To capture talent hoarding in the firm data, I construct a direct hoarding measure based on managers’ systematic suppression of worker visibility using two worker ratings that managers conduct. The ratings differ in the extent of their visibility outside of a worker’s team, meaning that managers can suppress worker visibility by lowering the public rating relative to the private rating. Such suppression of visibility is cited as a key form of talent hoarding in both managers’ and workers’ survey responses. Testing predictions from my conceptual framework based on three proxy measures for managerial hoarding incentives, I demonstrate that managers are indeed more likely to engage in talent hoarding if they face greater incentives to hoard. Additional evidence from alternative measures of talent hoarding based on managers restricting access to high-visibility trainings and managers’ explicit hoarding statements from the survey further corroborates this finding. I use quasi-random variation in workers’ exposure to talent hoarding based on manager rotations to demonstrate that managerial talent hoarding has meaningful impacts on worker outcomes. When a manager learns that they will move to a new position on a different team, they no longer have an incentive to hoard workers on their current team. For workers whose manager will soon rotate, this creates a temporary window of time during which they are not subject to talent hoarding. Consistent with the cessation of talent hoarding, I find that manager rotations increase worker applications by 78%. To support the interpretation of rotation effects as being driven by talent hoarding, I test predictions from the conceptual framework, which identifies specific subgroups of workers who should face higher levels of hoarding and therefore exhibit larger responses to manager rotations. Coupled with a series of tests for potentially confounding mechanisms, these empirical exercises support my interpretation based on talent hoarding and motivate the use of manager rotations to measure the impacts of talent hoarding on worker outcomes. Talent hoarding has important consequences for workers’ careers and the firm’s ability to allocate talent. To quantify how successful deterred applications would have been, I instrument for workers’ applications with manager rotations. Marginal applicants (identified via standard IV assumptions), who would not have applied in absence of a manager rotation and only apply because their manager rotates, face a 49.1% likelihood of landing a new position. This large and positive hiring likelihood suggests that many deterred applicants would have been successful and is particularly striking given that the average hiring likelihood is 27.6%. The positive selection of marginal applicants is further corroborated by a complier analysis (Abadie, 2003), which implies that talent hoarding has negative impacts on the composition of the applicant pool. Using workers’ performance ratings, I present evidence suggesting that a substantial share of deterred applicants would have been highly productive in their new positions. Together, my findings suggest that in the presence of talent hoarding, firms may forgo productivity gains from hoarded workers who are not assigned to positions in which they would have been productive. Talent hoarding may also exacerbate between-group inequality at the firm. A survey of the firm’s employees suggests that male and female workers may react differently to talent hoarding. In line with the literature on gender differences in preferences (Bertrand, 2011), the survey finds that women in the firm are 22% more likely to place a high value on preserving a good relationship with their manager and are 26% more likely to rely on managers’ career guidance when making career decisions. The effects of manager rotations corroborate that talent hoarding has differential effects by gender, affecting women at a higher part of the quality distribution compared to men. In particular, marginal female applicants are more qualified in terms of their educational qualifications and past performance. They are also more positively selected with respect to their hiring probability for higher-level positions. When comparing potential outcomes for marginal applicants, I find that both men and women experience higher earnings if they choose to apply in the absence of talent hoarding; however, the larger gains realized by women lead to a reduction in the pay gap of 86%. A number of factors suggest that talent hoarding is very likely to manifest similarly in other organizations. The firm I study is similar to other large firms in Germany both in terms of the characteristics of its workforce and its internal organization, in that it is standard that managers are tasked with identifying talented workers, but are neither monitored nor rewarded in doing so (hkp, 2021). Companies across the world report that talent hoarding is commonplace, creates barriers to talent allocation, and occurs through many of the same managerial behaviors that are documented in this study (i4cp, 2016, KornFerry, 2015, Matuson, 2015, Sullivan, 2017). This study contributes to two broad strands of research on organizations. Prior theoretical research has hypothesized that managers engage in self-interested behavior (Holmstrom and Tirole, 1989), largely focusing on managers’ misaligned incentives in the context of biased performance evaluations (Milgrom and Roberts, 1988, Prendergast and Topel, 1996, Fairburn and Malcomson, 2001). While research studying internal labor markets has documented the importance of incentive provision in organizations (Gibbons and Waldman, 1999, Prendergast, 1999), little attention has been paid to how managers’ incentive problems may affect the efficiency of job assignments. One notable exception is theoretical work by Friebel and Raith (2022) who show that different organizational designs may change managers’ incentives to train subordinates and accurately represent their abilities. My study provides the first empirical demonstration of a costly moral hazard problem among managers that likely affects the efficient allocation of talent within organizations. Second, a large empirical literature in economics studies the impacts of managers on firm outcomes. The majority of this literature has focused on upper management, and in particular on documenting the impacts of CEOs on firm performance (Bertrand and Schoar, 2003, Bennedsen et al., 2007, Malmendier and Tate, 2009, Bennedsen et al., 2020). An emerging body of work uses detailed data on managers and workers to show that even managers at lower levels of the firm hierarchy have large impacts on worker outcomes, including worker productivity (Lazear et al., 2015, Frederiksen et al., 2020, Fenizia, 2022), turnover (Hoffman and Tadelis, 2021), and career progression (Kunze and Miller, 2017, Cullen and Perez-Truglia, 2023, Benson et al., 2021). This study adds to these findings by uncovering talent hoarding as an important mechanism that influences managers’ impacts on firms and workers. By demonstrating that talent hoarding has meaningful impacts on career progression, this study also contributes to both theoretical and empirical work seeking to understand the dynamics of internal labor markets (Waldman, 1984, Milgrom and Oster, 1987, Baker et al., 1994, Benson et al., 2019, Huitfeldt et al., 2023). The rest of the paper proceeds as follows. Section 2 provides anecdotal evidence and introduces a simple framework that offers a formal definition of talent hoarding. Section 3 describes the institutional setting and data. Section 4 focuses on managers and provides direct evidence on the hoarding incentives they face and the actions that follow. Section 5 uses quasi-random variation centered around manager rotations to document the impact of managerial talent hoarding on workers. Section 6 presents evidence on the consequences of talent hoarding for workers’ careers and the firm’s ability to allocate talent. Section 7 concludes.
@@ 2 Background and Conceptual Framework
This section presents anecdotal evidence of how talent hoarding occurs in practice. I conduct a large-scale survey at the firm I study, which illustrates the reasons why and the actions through which managers hoard talent. Building on this intuition, a simple conceptual framework offers a formal definition of talent hoarding as well as a set of predictions that guides my empirical analysis. What is talent hoarding? A 2017 Wall Street Journal article describes talent hoarding as “manager’s natural tendency to hold on to top performers instead of working to promote them or transfer them to other areas of the company." (Lublin, 2017). Media outlets also describe how managers hoard talent. A 2015 Forbes article observes that managers who hoard talent “never recommend…people for a promotion in another department." (Matuson, 2015). The industry publication Talent Management & HR writes in 2017 that “hoarding managers, in order to reduce the internal visibility of their top team members, may purposely restrict them from serving on task forces and outside-of-function committees." (Sullivan, 2017). Other hoarding strategies are described as underrating potential for higher-level positions or threatening workers who try to leave the team. Talent hoarding is widespread and occurs in a variety of settings. In a survey of 665 global organizations, covering both the private and public sector, half of organizations report that their managers hoard talent by discouraging worker mobility (i4cp, 2016). In Germany, 83% of the top publicly listed companies cite talent hoarding as a crucial friction in their organization (hkp, 2021). Managers’ talent hoarding actions seem consequential. In one-third of US firms, workers feel the need to keep applications secret from their managers out of fear of retaliation (KornFerry, 2015).

### v3

@@
Talent Hoarding in Organizations
Ingrid Haegele [^1] Ludwig-Maximilians University; ingrid.haegele@econ.lmu.de. I thank David Card, Patrick Kline, and Christopher Walters for their advice and encouragement. This paper has benefited tremendously from comments and suggestions from Sydnee Caldwell, Florian Englmaier, Robert Gibbons, Lisa Kahn, Supreet Kaur, Danielle Li, Conrad Miller, Evan Rose, Jesse Rothstein, Steve Tadelis, Dmitry Taubinsky, Ricardo Perez-Truglia, Fabian Waldinger, Francis Wong, and numerous seminar and conference participants. I thank Leon Jacob and the employees of my institutional partner for their feedback and support. This work was supported by the Alfred P. Sloan Foundation Pre-doctoral Fellowship on the Economics of an Aging Workforce awarded through the NBER, IRLE Student Research Grant, O-LAB Labor Science Initiative Grant, Strandberg Grant for Gender in Economic Research, UC Berkeley Matrix Dissertation Fellowship, and Washington Center for Equitable Growth Doctoral Grant.
(July 2025)
Most organizations rely on managers to identify talented workers. However, managers who are evaluated on team performance have an incentive to hoard workers. This study provides the first empirical evidence of talent hoarding using personnel records and survey evidence from a large firm. Talent hoarding is self-reported by three-fourths of managers, is detectable in manager ratings of worker talent, and occurs more frequently under stronger hoarding incentives, proxied by performance-related pay, team size, and talent visibility. Using quasi-random exposure to talent hoarding, I show that hoarding deters internal job applications, inhibiting career progression and altering talent allocation in the firm.
@@ 1 Introduction
Firms must continually decide how to allocate workers to jobs, a process which has critical implications for productivity (Rosen, 1982, Holmstrom and Tirole, 1989). Because it is difficult to perfectly observe worker ability, most firms rely on managers to identify talented workers who can be promoted to higher-level positions. However, when a talented worker leaves their team for a new position, team performance suffers. Since managers are rewarded based on team performance and firms cannot perfectly monitor manager actions, the conflicting interests of manager and firm create the potential for moral hazard (Holmstrom, 1979). A growing body of evidence documents that workers in high-level positions have large impacts on firm performance (Bloom and Van Reenen, 2007, Lazear et al., 2015), implying that managers may create significant efficiency costs if they hoard talented workers rather than promote them.
@@
Ample anecdotal and survey evidence points to widespread talent hoarding in organizations. In a global survey, half of organizations report that managers hoard talent by discouraging worker mobility (i4cp, 2016). A US-based survey finds that workers in one-third of firms feel the need to keep internal applications secret from their managers out of fear of retaliation (KornFerry, 2015). In Germany, 83% of the top publicly listed companies cite managerial talent hoarding as a key friction in their organization (hkp, 2021). [^2] Talent hoarding also occurs in science and academia. Zuckerman (2021) documents how Katalin Karikó, a seminal developer of mRNA vaccines and 2023 Nobel prize winner, experienced talent hoarding when she decided to leave her lab for a new position. Her advisor Robert Suhadolnik vowed “to do whatever he could to stop his protégée from leaving…he made it clear she had two career choices. ‘You can work in my lab or go home,’ he told her. Suhadolnik followed through on this threat, telling a local immigration office that she was living in the country illegally and should be deported.” Despite the apparent prevalence of talent hoarding and its likely detrimental consequences, very little empirical evidence on talent hoarding exists in economics. Studying talent hoarding empirically is challenging. Managers often hoard talent through hidden actions that are difficult to observe, even in rich datasets. Furthermore, identifying the causal impacts of talent hoarding requires plausibly exogenous variation in hoarding.
This study provides the first empirical evidence on talent hoarding in organizations. I combine rich personnel records from a large manufacturing firm with survey evidence from both managers and workers in the firm. I document widespread managerial incentives to hoard talent: the majority of managers perceive a conflict of interest with respect to developing talent, because more developed workers are more likely to leave the team. Talent hoarding is reported by 75% of managers. Using a direct measure of talent hoarding derived from the firm’s personnel records, I test predictions from my conceptual framework and confirm that stronger hoarding incentives lead to more talent hoarding. I leverage quasi-random variation in workers’ exposure to talent hoarding to demonstrate the impacts of talent hoarding on workers’ career outcomes. Hoarding significantly deters workers’ applications for internal positions, many of which would have been successful. Because talent hoarding has a larger impact on high-quality workers, hoarding reduces the quality of the applicant pool, thus restricting the firm’s ability to effectively allocate talent within the organization.
I develop a simple conceptual framework that formalizes the misaligned incentives that give rise to talent hoarding. Since managers are both tasked with identifying productive workers and are rewarded based on team productivity, managers have an incentive to hoard talent by preventing workers from seeking new positions outside the team. This framework yields empirical predictions for both manager and worker behavior. For example, when gains to hoarding are larger or worker departure from the team is costlier, managers engage in more hoarding. High-productivity workers who experience more hoarding are more deterred from applying, which negatively impacts internal applicant pools and has the potential to distort the allocation of talent within the firm.
To empirically test for talent hoarding, I collect a unique combination of personnel records, internal job application data, and new survey evidence from a large manufacturing firm that employs over 200,000 workers. I focus my analysis on the firm’s largest internal labor market, consisting of over 30,000 white-collar and management employees in Germany. I demonstrate that the firm is similar to other large firms in terms of its workforce and organizational design. As in many other large firms, employees who want to switch to a new position or to be promoted are required to apply for the position. Since most workers in the same team work in similar positions, workers who seek different positions, including promotions, typically need to apply to jobs outside the team.
The data I assemble offer three key advantages. First, collecting new survey evidence from managers in the firm allows me to directly document that managers face misaligned incentives and act accordingly. Second, the firm’s personnel records allow me to construct a measure of talent hoarding using the difference between managers’ private and public ratings of worker talent. Without such data, directly measuring managerial talent hoarding is empirically challenging because hoarding by definition involves hidden actions. Third, by combining personnel records with the universe of application and hiring decisions at the firm, I am able to document the extent to which talent hoarding deters workers’ high-stakes applications that would otherwise have been successful.
Survey responses from 62% of managers in the firm reveal that incentives to hoard talent are widespread. The majority of managers (55%) explicitly acknowledge that supporting workers’ career development entails a conflict of interest because more developed workers are more likely to leave the team. This conflict of interest relates to the misaligned incentives managers face: while almost all managers (96%) believe they would have a large impact on workers’ careers, only 36% perceive this impact to be valued by the firm. A further disincentive to develop talent is the risk of workers leaving the team: 68% of managers state that they would be more likely to support workers’ career development if it would be easier to replace lost talent. The survey demonstrates that managerial talent hoarding actions, which are reported by 75% of managers, are a common response to these prevailing managerial incentives.
Managers’ revealed actions provide further evidence that managers purposefully hoard talent. To capture talent hoarding in the firm data, I construct a direct hoarding measure based on managers’ systematic suppression of worker visibility using the difference between the private performance ratings and public potential ratings that managers assign to workers. The ratings differ in the extent of their visibility outside of a worker’s team, meaning that managers can suppress worker visibility by lowering the public rating relative to the private rating. Such suppression of visibility is cited as a key form of talent hoarding in both managers’ and workers’ survey responses. Testing predictions from my conceptual framework based on three proxy measures for managerial hoarding incentives (i.e., performance-related pay, team size, and talent visibility), I demonstrate that managers are indeed more likely to engage in talent hoarding if they face greater incentives to hoard. Additional evidence from alternative measures of talent hoarding based on managers restricting access to high-visibility trainings and managers’ explicit hoarding statements from the survey further corroborates this finding.
These manager-level findings establish a clear link between managerial incentives and hoarding actions. I then examine whether these behaviors affect workers. To test whether managerial hoarding has meaningful consequences for worker outcomes, I use quasi-random variation in workers’ exposure to talent hoarding based on manager rotations. For workers whose manager learns that they will move to a new position on a different team, this creates a temporary window of time during which their manager does not have an incentive to hoard talent. Consistent with the cessation of talent hoarding, I find that manager rotations increase worker applications by 78%. To support the interpretation of rotation effects as being driven by talent hoarding, I test predictions from the conceptual framework, which predicts that workers with high productivity, those whose replacement is costlier, and those working under managers with lower utility costs of hoarding should exhibit larger responses to manager rotations. Coupled with a series of tests for potentially confounding mechanisms, such as worker-manager specific match effects and finite project timelines, these empirical exercises support my interpretation based on talent hoarding and motivate the use of manager rotations to measure the impacts of talent hoarding on worker outcomes.
Talent hoarding has important consequences for workers’ careers and the firm’s ability to allocate talent. To quantify how successful deterred applications would have been, I instrument for workers’ applications with manager rotations. Marginal applicants (identified via standard IV assumptions), who would not have applied in absence of a manager rotation and only apply because their manager rotates, face a 49.1% likelihood of landing a new position. This large and positive hiring likelihood suggests that many deterred applicants would have been successful and is particularly striking given that the average hiring likelihood is 27.6%. The positive selection of marginal applicants is further corroborated by a complier analysis (Abadie, 2003), which implies that talent hoarding has negative impacts on the composition of the applicant pool. Using workers’ performance ratings, I present evidence suggesting that a substantial share of deterred applicants would have been highly productive in their new positions. Together, my findings suggest that in the presence of talent hoarding, firms may forgo productivity gains from hoarded workers who are not assigned to positions in which they would have been productive.
Talent hoarding may also exacerbate between-group inequality at the firm. A survey of the firm’s employees suggests that male and female workers may react differently to talent hoarding. In line with the literature on gender differences in preferences (Bertrand, 2011), the survey finds that women in the firm are 22% more likely to place a high value on preserving a good relationship with their manager and are 26% more likely to rely on managers’ career guidance when making career decisions. The effects of manager rotations corroborate that talent hoarding has differential effects by gender, affecting women at a higher part of the quality distribution compared to men. In particular, marginal female applicants are more qualified in terms of their educational qualifications and past performance. They are also more positively selected with respect to their hiring probability for higher-level positions. When comparing potential outcomes for marginal applicants, I find that both men and women experience higher earnings if they choose to apply in the absence of talent hoarding; however, the larger gains realized by women lead to a reduction in the pay gap of 86%.
A number of factors suggest that talent hoarding is very likely to manifest similarly in other organizations. The firm I study is similar to other large firms in Germany both in terms of the characteristics of its workforce and its internal organization, in that it is standard that managers are tasked with identifying talented workers, but are neither monitored nor rewarded in doing so (hkp, 2021). Companies across the world report that talent hoarding is commonplace, creates barriers to talent allocation, and occurs through many of the same managerial behaviors that are documented in this study (i4cp, 2016, KornFerry, 2015, Matuson, 2015, Sullivan, 2017).
This study contributes to two broad strands of research on organizations. Prior theoretical research has hypothesized that managers engage in self-interested behavior (Holmstrom and Tirole, 1989), largely focusing on managers’ misaligned incentives in the context of biased performance evaluations (Milgrom and Roberts, 1988, Prendergast and Topel, 1996, Fairburn and Malcomson, 2001). While research studying internal labor markets has documented the importance of incentive provision in organizations (Gibbons and Waldman, 1999, Prendergast, 1999), little attention has been paid to how managers’ incentive problems may affect the efficiency of job assignments. One notable exception is theoretical work by Friebel and Raith (2022) who show that different organizational designs may change managers’ incentives to train subordinates and accurately represent their abilities. My study provides the first empirical demonstration of a costly moral hazard problem among managers that likely affects the efficient allocation of talent within organizations.
Second, a large empirical literature in economics studies the impacts of managers on firm outcomes. The majority of this literature has focused on upper management, and in particular on documenting the impacts of CEOs on firm performance (Bertrand and Schoar, 2003, Bennedsen et al., 2007, Malmendier and Tate, 2009, Bennedsen et al., 2020). An emerging body of work uses detailed data on managers and workers to show that even managers at lower levels of the firm hierarchy have large impacts on worker outcomes, including worker productivity (Lazear et al., 2015, Frederiksen et al., 2020, Fenizia, 2022), turnover (Hoffman and Tadelis, 2021), and career progression (Kunze and Miller, 2017, Cullen and Perez-Truglia, 2023, Benson et al., 2021). This study adds to these findings by uncovering talent hoarding as an important mechanism that influences managers’ impacts on firms and workers. By demonstrating that talent hoarding has meaningful impacts on career progression, this study also contributes to both theoretical and empirical work seeking to understand the dynamics of internal labor markets (Waldman, 1984, Milgrom and Oster, 1987, Baker et al., 1994, Benson et al., 2019, Huitfeldt et al., 2023).
The rest of the paper proceeds as follows. Section 2 provides anecdotal evidence and introduces a simple framework that offers a formal definition of talent hoarding. Section 3 describes the institutional setting and data. Section 4 focuses on managers and provides direct evidence on the hoarding incentives they face and the actions that follow. Section 5 uses quasi-random variation centered around manager rotations to document the impact of managerial talent hoarding on workers. Section 6 presents evidence on the consequences of talent hoarding for workers’ careers and the firm’s ability to allocate talent. Section 7 concludes.
@@ 2 Background and Conceptual Framework
This section presents anecdotal evidence of how talent hoarding occurs in practice. I conduct a large-scale survey at the firm I study, which illustrates the reasons why and the actions through which managers hoard talent. Building on this intuition, a simple conceptual framework offers a formal definition of talent hoarding as well as a set of predictions that guides my empirical analysis.
What is talent hoarding? A 2017 Wall Street Journal article describes talent hoarding as “manager’s natural tendency to hold on to top performers instead of working to promote them or transfer them to other areas of the company.” (Lublin, 2017). Media outlets also describe how managers hoard talent. A 2015 Forbes article observes that managers who hoard talent “never recommend…people for a promotion in another department.” (Matuson, 2015). The industry publication Talent Management & HR writes in 2017 that “hoarding managers, in order to reduce the internal visibility of their top team members, may purposely restrict them from serving on task forces and outside-of-function committees” (Sullivan, 2017). Other hoarding strategies are described as underrating potential for higher-level positions or threatening workers who try to leave the team.
Talent hoarding is widespread and occurs in a variety of settings. In a survey of 665 global organizations, covering both the private and public sector, half of organizations report that their managers hoard talent by discouraging worker mobility (i4cp, 2016). In Germany, 83% of the top publicly listed companies cite talent hoarding as a crucial friction in their organization (hkp, 2021). Managers’ talent hoarding actions seem consequential. In one-third of US firms, workers feel the need to keep applications secret from their managers out of fear of retaliation (KornFerry, 2015).